It costs the johnsons $216 to have a 9 ft by 12 ft wooden floor refinished. At that rate,how muchwould it cost them to have a 12 ft by 16 ft wooden floor refinished?
Answer by rfer(16322)

(9*12)/216=(12*16)/x
108/216=192/x
108x=41472
x=41472/108
x=$384